OpenStudy (anonymous):

HELP! MEDALS WILL BE REWARDED AND I WILL FAN WHO EVER HELPS ME! A square table has a top that has an area of 11 square feet. To the nearest hundredth, what length of edging is needed to go around all edges of the tabletop and explain your answer? a.= 28.84 ft b.=13.27 ft c.=40.36 ft d.=13.97 ft

OpenStudy (anonymous):

do you know the formula for an area of a square

OpenStudy (anonymous):

never mind that doesn't apply here. all you need to do is find the perimeter of the square.

OpenStudy (anonymous):

I can't remember... wait is it this formula A=l*w?

OpenStudy (anonymous):

so is it 11+11+11+11=44?

OpenStudy (anonymous):

no a = length of one side squared

OpenStudy (anonymous):

so we need to find the length of one side , then multiply that by 4

OpenStudy (anonymous):

So do we multiply 11 by 4?

OpenStudy (anonymous):

no, 11 is the area.

OpenStudy (anonymous):

we need the length of one side of the square

OpenStudy (anonymous):

So do we use inverse operation and divide 11 by 4?

OpenStudy (anonymous):

no. 11 is the AREA. we need the measurement of the perimeter.

OpenStudy (anonymous):

we do this by doing the inverse formula, so 11=l^2. l is the length of one side

OpenStudy (anonymous):

I know 11 is the AREA! I just thought if the length times the width = 11. maybe if we divide 11 by 4 it will get the length.

OpenStudy (anonymous):

you're not using the right formula.

OpenStudy (anonymous):

 11sqaure feet

OpenStudy (anonymous):

the formula is l^2 not lxw

OpenStudy (anonymous):

Then help me use the right formula "plohrr".

OpenStudy (anonymous):

I just typed in the right formula above lol a = l^2

OpenStudy (anonymous):

now sub in the area for a, so it would be 11=l^2

OpenStudy (anonymous):

now we square both sides. so the square root of 11 and the square root of s^2

OpenStudy (anonymous):

I get where your going, but I'm not sure how your getting there. I know the area is 11 right. I also know the length is being squared but I don't know the length still?

OpenStudy (anonymous):

we're getting there

OpenStudy (anonymous):

How long is it going to take us because I'm going to bed at 12:00am.

OpenStudy (anonymous):

after we square both sides, we get 3.31 = l

OpenStudy (anonymous):

I don't get how you got that but at least I know the length is 3.31 now! :)

OpenStudy (anonymous):

this means each side is 3.31 long.

OpenStudy (anonymous):

multiply by 4 and get 13.27

OpenStudy (anonymous):

So 3.31 *4 = 13.27! THANKS!

OpenStudy (anonymous):

yep no problemo